# Video Understanding with Large Language Models: A Survey

### v3

@@ I Introduction
We live in a multimodal world where video has become the predominant form of media. With the rapid expansion of online video platforms and the growing prevalence of cameras in surveillance, entertainment, and autonomous driving, video content has risen to prominence as a highly engaging and rich medium, outshining traditional text and image-text combinations in both depth and appeal. This advancement has promoted an exponential surge in video production, with millions of videos being created daily. However, manually processing such a sheer volume of video content are labor-intensive and time-consuming. Consequently, there is a growing demand for tools to effectively manage, analyze, and process this abundance of video content. To meet that, video understanding methods and analysis technologies have emerged, leveraging intelligent analysis techniques to automatically recognize and interpret video content, thus significantly reducing the workload on human operators. Moreover, the continuous advancements in these methods enhance their task-solving capabilities, enabling them to handle a diverse range of video understanding tasks with increasing proficiency.
@@ I-A2 Early Neural Video Models
Compared with classical methods, deep learning methods for video understanding possess superior task-solving capabilities. DeepVideo [14] was the earliest method introducing a deep neural network, specifically a Convolutional Neural Network (CNN), for video understanding. However, the performance was not superior to the best handcrafted feature method due to the inadequate use of motion information. Two-stream networks [15] combined both CNN and IDT to capture the motion information to improve the performance, which verified the capability of deep neural networks for video understanding. To handle long-form video understanding, Long Short-Term Memory (LSTM) was adopted [16]. Temporal Segment Network (TSN) [17] was also designed for long-form video understanding by analyzing and aggregating video segments. Based on TSN, Fisher Vectors (FV) encoding [18], Bi-Linear encoding [19], and Vector of Locally Aggregated Descriptors (VLAD) [20] encoding were introduced [21]. These methods improved performance on the UCF-101 [22] and HMDB51 [23] datasets. Different from two-stream networks, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) [24]. Inflated 3D ConvNets (I3D) [25] utilizes the initialization and the architecture of 2D CNN, Inception [26], to gain a huge improvement on the UCF-101 and HMDB51 datasets. Subsequently, people began employing the Kinetics-400 (K-400) [27] and Something-Something [28] datasets to evaluate the model’s performance in more challenging scenarios. ResNet [29], ResNeXt [30], and SENet [31] were also adapted from 2D to 3D, resulting in the emergence of R3D [32], MFNet [33], and STC [34]. To improve the efficiency, the 3D network has been decomposed into 2D and 1D networks in various studies (e.g., S3D [35], ECO [36], P3D [37]). LTC [38], T3D [39], Non-local [40], and V4D [41] focus on long-form temporal modeling, while CSN [42], SlowFast [43], and X3D [44] tend to attain high efficiency. The introduction of Vision Transformers (ViT) [45] promotes a series of prominent models (e.g., TimeSformer [46], VidTr [47], ViViT [48], MViT [49]).
Figure 2: A comprehensive timeline depicting the development of video understanding methods with large language models (Vid-LLMs) over the recent months. This survey of Vid-LLMs is based on advancements up to the end of June 2024.
@@ I-A3 Self-supervised Video Pretraining
Transferability [50, 51] in self-supervised pretraining models [52] for video understanding allows them to generalize across diverse tasks with minimal additional labeling, overcoming the early deep learning models’ requirements for extensive task-specific data. VideoBERT [53] is an early attempt to perform video pre-training. Based on the bidirectional language model BERT [54], pertaining tasks are designed for self-supervised learning from video-text data. It tokenizes video features with hierarchical k-means. The pre-trained model can be fine-tuned to handle multiple downstream tasks, including action classification and video captioning. Following the ”pre-training” and ”fine-tuning” paradigm, a large number of studies on pre-trained models for video understanding, especially video-language models, have emerged. They either use different architectures (ActBERT [55], SpatiotemporalMAE [56], OmniMAE [57], VideoMAE [58], MotionMAE [59]) or pre-training and fine-tuning strategies (MaskFeat [60], VLM [61], ALPRO [62], All-in-One transformer [63], maskViT [64], CLIP-ViP [65], Singularity [66], LF-VILA [67], EMCL [68], HiTeA [69], CHAMPAGNE [70]).
@@ I-A4 Large Language Models for Video Understanding
Recently, large language models (LLMs) have advanced rapidly [71]. The emergence of large language models pre-trained on extensive datasets has introduced a novel in-context learning capability [72]. This allows them to handle various tasks using prompts without the need for fine-tuning. ChatGPT [73] is the first groundbreaking application built on this foundation. This includes capabilities like generating code and invoking tools or APIs of other models for their use. Many studies are exploring to use of LLMs like ChatGPT to call vision models APIs to solve the problems in the computer vision field, including Visual-ChatGPT [74]. The advent of instruct-tuning has further enhanced these models’ ability to respond effectively to user requests and perform specific tasks. LLMs integrated with video understanding capabilities offer the advantage of more sophisticated multimodal understanding, enabling them to process and interpret complex interactions between visual and textual data. Similar to their impact in Natural Language Processing (NLP) [75], these models act as more general-purpose task solvers, adept at handling a broader range of tasks by leveraging their extensive knowledge base and contextual understanding acquired from vast amounts of multimodal data. This allows them to not only understand visual content but also reason about it in a way that is more aligned with human-like understanding. Many works also explore using LLMs in video understanding tasks, namely, Vid-LLMs.
@@ I-C Survey Structure
This survey is structured as follows: Section II offers preliminaries for video understanding with LLMs, including a summary of various video understanding tasks that require handling different levels of granularity, their associated datasets, and evaluation metrics. The background of LLMs is also introduced in this section. In Section III, we delve into details of recent research leveraging LLMs for video understanding, presenting their unique approaches and impact in the field, where we divide these Vid-LLMs into three main categories, Video Analyzer $\times$ LLM and Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM; and five sub-categories, LLM as Summarizer/Manger/Text Decoder/Regressor/Hidden Layer, shown as Figure 4. This section also includes the training strategies of Vid-LLMs. Section IV adds more information about popular ways to evaluate Vid-LLMs, together with some benchmarks and performances of some Vid-LLMs on the most commonly-used benchmarks. Section V explores the application of Vid-LLMs across multiple significant fields. The survey concludes in Section VI by summarizing key findings and identifying unresolved challenges and potential areas for future research.
@@ II-A Video Understanding Tasks
Video understanding is a challenging domain that has inspired the creation of numerous innovative tasks to enhance models’ capabilities to interpret video content. The field has evolved from foundational tasks such as video classification and action recognition, which categorize videos into predefined categories. As technology has advanced, the scope of video understanding has expanded to include more complex tasks. For instance, video captioning tasks not only require the model to provide detailed descriptions of the video content but also necessitate logical and commonsense reasoning about the scenes depicted. Additionally, video question-answering tasks demand that the model not only understands the content but also integrates external information to provide accurate responses. These more intricate tasks illustrate a natural progression from simple categorization to deeper comprehension and reasoning, highlighting the increasing demand for models to achieve near-human levels of video interpretation. We summarize the main tasks in video understanding as follows.
Figure 3: This figure categorizes tasks in video understanding, delineating the granularity required and the language involvement necessary for models to perform these tasks effectively. This diagram excludes tasks that involve special modalities or specific, such as audio-visual video understanding and egocentric video understanding. Notably, the tasks presented could be unified into a question-answering paradigm, and all solved by generative large models, akin to recent advances in NLP.
@@ II-A1 Abstract Understanding Tasks
  • Video Classification & Action Recognition: Video classification and action recognition involve identifying and categorizing activities or events within a video sequence. Different datasets to solve these tasks include UCF-101 [22], HMDB51 [23], Hollywood [81], ActivityNet [82], Charades [83], Kinetics-400 [27], Kinetics-600 [84], Kinetics-700 [85], SomethingSomethingV2 [86], HACS [87], YouTube8M [88], and PortraitMode-400 [89]. Usually, Top-K accuracy is adopted as the main metric for these tasks.
  • Text-Video Retrieval: Text-video retrieval is a task that involves matching and retrieving relevant video clips based on textual descriptions. Datasets like Kinetic-GEB [90], MSRVTT [91], DiDeMo [92], YouCook2 [93], and Youku-mPLUG [94] are relevant to this task. The standard evaluation metric for this task is Recall at K (R@K), which measures the accuracy of the first K retrieved results.
  • Video-to-Text Summarization: Video-to-text summarization is a task that generates concise textual summaries from video content. This process extracts and interprets key visual and audio elements to produce coherent and informative summaries. ViTT [95], VideoXum [96], VideoInstruct-100K [97], and Instrcut-V2Xum [98] are datasets that related to the task. This task is often evaluated by metrics of BLEU, METEOR, CIDEr, and ROUGE-L.
  • Video Captioning: Video captioning is the task of generating descriptive and coherent textual captions for video content. It interprets visual and auditory information from video frames to produce accurate and contextually relevant descriptions. Notable datasets for this tasks are MSVD [99], MSR-VTT [100], TGIF [101], Charades [83], Charades-Ego [102], YouCook2 [93], Youku-mPLUG [94], VAST-27M [103], and VideoInstruct-100K [97]. This task is often evaluated by metrics of BLEU, METEOR, CIDEr, and ROUGE-L.
  • Video QA: Video QA (Question Answering) is about answering natural language questions based on the content of a video to analyze visual and auditory information, understand the context, and generate accurate responses. Datasets involved in the QA task are VCR [104], MSVD-QA [91], MSRVTT-QA [91], TGIF-QA [105], Pororo-QA [106], TVQA [107], ActivityNet-QA [108], and NExT-QA [109]. This task is evaluated using Top-1, Top-K accuracy.
@@ II-A2 Temporal Understanding Tasks
  • Video Summarization: Video Summarization aims at condensing a long video into a shorter version while preserving essential content and important information. This task is usually evaluated by F1-score, Spearman, and Kendall as metrics. Commonly-used datasets include SumMe [110], TVSum [111], Ads-1k [112], VideoXum [96], Instrcut-V2Xum [98].
  • Video Highlight Detection: Video highlight detection is the process of identifying and extracting the most important and interesting segments from a video. Commonly used datasets on this task are the YouTube Highlights dataset [113], the TV-Sum dataset [111], and the Videos Titles in the Wild (VTW) dataset [114].
  • Temporal Action/Event Localization: Temporal action/event localization involves identifying the precise temporal boundaries of actions or events within a video. By analyzing sequential frames, this task pinpoints when specific activities start and end. Datasets for Temporal Action/Event Localization involve THUMOS’14 [115], ActivityNet-1.3 [82], and UnAV-100 [116].
@@
  • Video Temporal Grounding: Video temporal grounding is the task of locating specific moments or intervals within a video that correspond to a given textual query. This process involves aligning linguistic descriptions with visual content, enabling precise identification of relevant segments for applications in video search and content analysis. Common benchmarks are Charades-STA [117], ViTT [95], DiDeMo [92], and PU-VALOR [118]. This task is often evaluated by the metrics of R1@0.5 and R1@0.7.
@@
  • Generic Event Boundary Detection: Generic event boundary detection involves identifying the points in a video where significant changes or transitions occur and marking the boundaries between different events or activities. Kinetics-GEBD [119] is one of the datasets for this task.
@@
  • Dense Video Captioning: Dense video captioning involves generating detailed and continuous textual descriptions for multiple events and actions occurring throughout a video. Evaluation metrics like BLEU, METEOR, CIDEr, and ROUGE-L are used to evaluate this task. Relevant datasets are ActivityNet Captions [120], VidChapters-7M [121], YouCook2 [93], and ViTT [95].
@@ II-A3 Spatiotemporal Understanding Tasks
  • Object Tracking: Object tracking involves continuously identifying and following the position of specific objects within a video over time. It aims to maintain accurate and consistent trajectories of objects despite challenges like occlusions, changes in appearance, and motion. Benchmarks like OTB [122], UAV [123], and VOT [124] are commonly used for this task.
  • ReID: Re-Identification (ReID) is the task of recognizing and matching individuals or objects across different video frames or camera views. Common datasets in ReID are Market-1501 [125], CUHK03 [126], MSMT17 [127], and DukeMTMC-reID [128].
  • Video Saliency Detection: Video saliency detection involves identifying the most visually important and attention-grabbing regions in a video. This task highlights areas that stand out due to factors like motion, contrast, and unique features. Relevant datasets to this task are DHF1K [129], Hollywood-[83], UCF-Sports [130], AVAD [131], Coutrot1 [132], Coutrot2 [133], ETMD [134], and SumMe [110].
  • Video Object Segmentation: Video object segmentation involves partitioning a video into segments that correspond to individual objects, accurately delineating their boundaries over time. YouTube-VOS [135] and DAVIS [136] are datasets related to this task.
  • Video Instance Segmentation: Video instance segmentation is the task of identifying, segmenting, and tracking each unique object instance within a video. YouTube-VIS [137] and Cityscapes-Seq [138] are benchmarks for this task.
  • Video Object Referring Segmentation: Video object referring segmentation involves segmenting specific objects in a video based on natural language descriptions. It identifies and isolates the referred objects accurately across frames, where MeViS [139] is the benchmark dataset for this task.
  • Spatiotemporal Grounding: Spatiotemporal grounding involves identifying and localizing specific objects or events within both spatial and temporal dimensions of a video based on a given query. Datasets like Vid-STG [140], HC-STVG [141], Ego4D-MQ and Ego4D-NLQ [142] are proposed to aid the training and testing for this task.

### v4

@@ I Introduction
We live in a multimodal world where video has become the predominant form of media. With the rapid expansion of online video platforms and the growing prevalence of cameras in surveillance, entertainment, and autonomous driving, video content has risen to prominence as a highly engaging and rich medium, outshining traditional text and image-text combinations in both depth and appeal. This advancement has promoted an exponential surge in video production, with millions of videos being created daily. However, manually processing such a sheer volume of video content is labor-intensive and time-consuming. Consequently, there is a growing demand for tools to effectively manage, analyze, and process this abundance of video content. To meet that, video understanding methods and analysis technologies have emerged, leveraging intelligent analysis techniques to automatically recognize and interpret video content, thus significantly reducing the workload on human operators. Moreover, the continuous advancements in these methods enhance their task-solving capabilities, enabling them to handle a diverse range of video understanding tasks with increasing proficiency.
@@ I-A2 Early Neural Video Models
Compared with classical methods, deep learning methods for video understanding possess superior task-solving capabilities. DeepVideo [14] and [15] were early methods introducing a deep neural network, specifically a Convolutional Neural Network (CNN), for video understanding. However, the performance was not superior to the best handcrafted feature method due to the inadequate use of motion information. Two-stream networks [16] combined both CNN and IDT to capture the motion information to improve the performance, which verified the capability of deep neural networks for video understanding. To handle long-form video understanding, Long Short-Term Memory (LSTM) was adopted [17]. Temporal Segment Network (TSN) [18] was also designed for long-form video understanding by analyzing and aggregating video segments. Besides TSN, Fisher Vectors (FV) encoding [19], Bi-Linear encoding [20], and Vector of Locally Aggregated Descriptors (VLAD) [21] encoding were introduced [22]. These methods improved performance on the UCF-101 [23] and HMDB51 [24] datasets. Unlike two-stream networks, 3D networks started another branch by introducing 3D CNN to video understanding (C3D) [25]. Inflated 3D ConvNets (I3D) [26] utilizes the initialization and the architecture of 2D CNN, Inception [27], to gain a huge improvement on the UCF-101 and HMDB51 datasets. Subsequently, people began employing the Kinetics-400 (K-400) [28] and Something-Something [29] datasets to evaluate the model’s performance in more challenging scenarios. ResNet [30], ResNeXt [31], and SENet [32] were also adapted from 2D to 3D, resulting in the emergence of R3D [33], MFNet [34], and STC [35]. To improve the efficiency, the 3D convolution has been decomposed into cascade 2D and 1D convolution in various studies (e.g., S3D [36], ECO [37], P3D [38]). LTC [39], T3D [40], Non-local [41], and V4D [42] focus on long-form temporal modeling, while CSN [43], SlowFast [44], and X3D [45] tend to attain high efficiency. The introduction of Vision Transformers (ViT) [46] promotes a series of prominent models (e.g., TimeSformer [47], VidTr [48], ViViT [49], MViT [50]).
Figure 2: A comprehensive timeline depicting the development of video understanding methods with large language models (Vid-LLMs). This survey is based on advancements up to the end of June 2024.
@@ I-A3 Self-supervised Video Pretraining
Transferability [51, 52] in self-supervised pretraining models [53] for video understanding allows them to generalize across diverse tasks with minimal additional labeling, overcoming the early deep learning models’ requirements for extensive task-specific data. VideoBERT [54] is an early attempt to perform video pretraining. Based on the bidirectional language model BERT [55], pertaining tasks are designed for self-supervised learning from video-text data. It tokenizes video features with hierarchical k-means. The pretrained model can be fine-tuned to handle multiple downstream tasks, including action classification and video captioning. Following the “pretraining-finetuning” paradigm, many studies on pretrained models for video understanding, especially video-language models, have emerged. They either use different architectures (ActBERT [56], SpatiotemporalMAE [57], OmniMAE [58], VideoMAE [59], MotionMAE [60]) or training strategies (MaskFeat [61], VLM [62], ALPRO [63], All-in-One transformer [64], MaskViT [65], CLIP-ViP [66], Singularity [67], LF-VILA [68], EMCL [69], HiTeA [70], CHAMPAGNE [71]).
@@ I-A4 Large Language Models for Video Understanding
Recently, large language models (LLMs) have advanced rapidly [72]. The emergence of large language models pretrained on extensive datasets has introduced a novel in-context learning capability [73]. This allows them to handle various tasks using prompts without the need for fine-tuning. ChatGPT [74] is the first groundbreaking application built on this foundation. This includes capabilities like generating code and invoking tools or APIs of other models for their use. Many studies are exploring using LLMs like ChatGPT to call vision models APIs to solve the problems in the computer vision field, including Visual-ChatGPT [75]. The advent of instruct-tuning has further enhanced these models’ ability to respond effectively to user requests and perform specific tasks [76]. LLMs integrated with video understanding capabilities offer the advantage of more sophisticated multimodal understanding, enabling them to process and interpret complex interactions between visual and textual data. Similar to their impact in Natural Language Processing (NLP) [77], these models act as more general-purpose task solvers, adept at handling a broader range of tasks by leveraging their extensive knowledge base and contextual understanding acquired from vast amounts of multimodal data. This allows them to not only understand visual content but also reason about it in a way that is more aligned with human-like understanding. Many works also explore using LLMs in video understanding tasks, namely, Vid-LLMs.
@@ I-C Survey Structure
This survey is structured as follows: Section II offers preliminaries for video understanding with LLMs, including a summary of various video understanding tasks that require handling different levels of granularity, their associated datasets, and evaluation metrics. The background of LLMs is also introduced in this section. In Section III, we delve into details of recent research leveraging LLMs for video understanding, presenting their unique approaches and impact in the field, where we divide these Vid-LLMs into three main categories, Video Analyzer $\times$ LLM and Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM; and five sub-categories, LLM as Summarizer/Manager/Text Decoder/Regressor/Hidden Layer, shown as Figure 5. This section also includes the training strategies of Vid-LLMs. Section IV adds more information about popular ways to evaluate Vid-LLMs, together with some benchmarks and performances of some Vid-LLMs on the most commonly used benchmarks. Section V explores the application of Vid-LLMs across multiple significant fields. The survey concludes in Section VI by summarizing key findings and identifying unresolved challenges and potential areas for future research.
@@ II-A Video Understanding Tasks
Video understanding is a fundamental yet challenging task that has inspired the emergence of numerous tasks in a similar discipline aiming at interpreting complicated video content. The pioneering work for video understanding includes video classification and action recognition approaches, which classify videos into class labels and action categories, respectively. With the development of visual foundation models and expanding public datasets, current video understanding approaches can capture, analyze, and reason for more complicated video content. For instance, video captioning, as a specific task of video understanding, not only requires the model to generate detailed descriptions of the video content, but the generated video captions should be logical and follow commonsense about the scenes depicted. Additionally, the Video Question-Answering (VQA) task requires that the model understand the content and refer to external information to provide an accurate answer. The development path of video understanding from simple classification to natural language comprehension and reasoning highlights a clear trend of the video understanding model towards near-human levels of video interpretation capability. We summarize the main tasks in video understanding as follows.
Figure 3: This figure categorizes tasks in video understanding, delineating the granularity required and the language involvement necessary for models to perform these tasks effectively. This diagram excludes tasks involving special modalities or specific, such as audio-visual and egocentric video understanding. Notably, the tasks presented could be unified into a question-answering paradigm, and all solved by generative large models, akin to recent advances in NLP.
@@ II-A1 Abstract Understanding Tasks
  • Video Classification & Action Recognition: Video classification and action recognition classify videos based on class labels or activities and events categories within a video sequence. Datasets specifically introduced for these tasks include UCF-101 [23], HMDB51 [24], Hollywood [83], ActivityNet [84], Charades [85], Kinetics-400 [28], Kinetics-600 [86], Kinetics-700 [87], SomethingSomethingV2 [88], HACS [89], YouTube8M [90], and PortraitMode-400 [91]. Usually, Top-K accuracy is adopted as the main metric for these tasks.
  • Text-Video Retrieval: Text-video retrieval task matches and retrieves relevant video clips based on the similarity between video clips and the input textual descriptions. Datasets like Kinetic-GEB [92], MSRVTT [93], DiDeMo [94], YouCook2 [95], and Youku-mPLUG [96] are relevant to this task. The standard evaluation metric for this task is Recall at K (R@K), which measures the accuracy of the first K retrieved results.
  • Video-to-Text Summarization: Video-to-text summarization is a task that generates concise textual summaries of videos. The video summarization approaches are trained to extract and interpret key visual and audio content to produce coherent and informative summaries. ViTT [97], VideoXum [98], VideoInstruct-100K [99], and Instrcut-V2Xum [100] are datasets that related to the task. Metrics of BLEU, METEOR, CIDEr, and ROUGE-L often evaluate this task.
  • Video Captioning: Video captioning generates descriptive and coherent textual captions of given videos. The video caption models usually use visual and auditory information from video to produce accurate and contextually relevant descriptions. Notable datasets for this tasks are MSVD [101], MSR-VTT [102], TGIF [103], Charades [85], Charades-Ego [104], YouCook2 [95], Youku-mPLUG [96], VAST-27M [105], and VideoInstruct-100K [99]. This task is often evaluated by metrics of BLEU, METEOR, CIDEr, and ROUGE-L.
  • Video QA: Video Question-Answering (VQA) aims to answer textual questions based on a given video, where the model analyzes visual and auditory information, understands the context, and eventually generates accurate responses. Datasets involved in the QA task are VCR [106], MSVD-QA [93], MSRVTT-QA [93], TGIF-QA [107], Pororo-QA [108], TVQA [109], ActivityNet-QA [110], and NExT-QA [111]. This task is evaluated using Top-1, Top-K accuracy.
@@ II-A2 Temporal Understanding Tasks
  • Video Summarization: Video Summarization aims at condensing a long video into a shorter version while preserving essential content and important information. F1-score, Spearman, and Kendall usually evaluate this task as metrics. Commonly-used datasets include SumMe [112], TVSum [113], Ads-1k [114], VideoXum [98], Instrcut-V2Xum [100].
  • Video Highlight Detection: Video highlight detection aims at identifying and extracting the most important and interesting segments from a video. Commonly used datasets on this task include the YouTube Highlights [115], the TV-Sum [113], and the Videos Titles in the Wild (VTW) [116].
  • Temporal Action/Event Localization: Temporal action/event localization task aims at identifying the precise temporal segments of actions or events within a video. By analyzing sequential frames, models trained for this task must indicate when specific activities start and end. Datasets for Temporal Action/Event Localization involve THUMOS’14 [117], ActivityNet-1.3 [84], and UnAV-100 [118].
@@
  • Video Temporal Grounding: Video temporal grounding is the task of locating specific moments or intervals within a video that correspond to a given textual query. This process involves aligning linguistic descriptions with visual content, enabling precise identification of relevant segments for applications in video search and content analysis. Common benchmarks are Charades-STA [119], ViTT [97], DiDeMo [94], and PU-VALOR [120]. The metrics of R1@0.5 and R1@0.7 often evaluate this task.
@@
  • Generic Event Boundary Detection: Generic event boundary detection involves identifying certain frames in a video where significant changes or transitions occur and splitting videos based on different events or activities. Kinetics-GEBD [121] is a widely-used dataset for this task.
@@
  • Dense Video Captioning: Dense video captioning aims at generating detailed and continuous textual descriptions for multiple events and actions occurring throughout a video. Evaluation metrics like BLEU, METEOR, CIDEr, and ROUGE-L are used to evaluate this task. Relevant datasets are ActivityNet Captions [122], VidChapters-7M [123], YouCook2 [95], and ViTT [97].
@@ II-A3 Spatiotemporal Understanding Tasks
  • Object Tracking: Object tracking aims at continuously identifying and following the position of specific objects within a video over time. A good object tracking model should maintain accurate and consistent tracking trajectories of objects, even for videos with occlusions, changes in appearance, and motions. Benchmarks like OTB [124], UAV [125], and VOT [126] are commonly used for this task.
  • Re-Identification: Re-Identification (ReID) is the task of recognizing and matching individuals or objects across different video frames or camera views. Common datasets in ReID are Market-1501 [127], CUHK03 [128], MSMT17 [129], and DukeMTMC-reID [130].
  • Video Saliency Detection: Video saliency detection aims at identifying the most visually important and attention-grabbing regions in a video. This task highlights areas that stand out due to factors like motion, contrast, and unique features. Relevant datasets to this task are DHF1K [131], Hollywood-[85], UCF-Sports [132], AVAD [133], Coutrot1 [134], Coutrot2 [135], ETMD [136], and SumMe [112].
  • Video Object Segmentation: Video object segmentation aims at partitioning a video into segments that correspond to individual objects, accurately delineating their boundaries over time. YouTube-VOS [137] and DAVIS [138] are datasets related to this task.
  • Video Instance Segmentation: Video instance segmentation is the task of identifying, segmenting, and tracking each unique object instance within a video. YouTube-VIS [139] and Cityscapes-Seq [140] are two common benchmarks for this task.
  • Video Object Referring Segmentation: Video object referring segmentation involves segmenting specific objects in a video based on natural language descriptions. It identifies and isolates the referred objects accurately across frames, where MeViS [141] is a common benchmark for this task.
  • Spatiotemporal Grounding: Spatiotemporal grounding aims to identify and localize specific objects or events within a video’s spatial and temporal dimensions based on a given query. Datasets like Vid-STG [142], HC-STVG [143], Ego4D-MQ and Ego4D-NLQ [144] are proposed to aid the training and testing for this task.